### v1

# Separable models for dynamic signed networks

Abstract
Signed networks capture the polarity of relationships between nodes, providing valuable insights into complex systems where both supportive and antagonistic interactions play a critical role in shaping the network’s dynamics. We propose a separable temporal generative framework based on multi-layer exponential random graph models, characterised by the assumption of conditional independence between the sign and interaction effects. This structure preserves the flexibly and explanatory power inherent in the binary network specification while adhering to consistent balance theory assumptions. Using a fully probabilistic Bayesian paradigm, we infer the doubly intractable posterior distribution of model parameters via an adaptive Metropolis-Hastings approximate exchange algorithm. We illustrate the interpretability of our model by analysing signed relations among U.S. Senators during Ronald Reagan’s second term (1985–1989). Specifically, we aim to understand whether these relations are consistent and balanced or reflect patterns of supportive or antagonistic alliances.
Dynamic networks,
Exponential random graph models,
Bayesian inference,
keywords:
[…]
and
1 Introduction
Network data characterise the connectivity structures of complex systems, which often involve a delicate balance between two types of interactions. In biological systems, for example, these interactions may be reflected in co-expression similarities and dissimilarities between genes (Mason et al., 2009) that are explained by the sign of the correlation of their expression profiles, while in social or political networks, they manifest as alliances and rivalries between individuals (Fontan and Altafini, 2021).
Signed networks, which consist of positive and negative edges, provide a powerful tool for capturing the nuanced relationships within these systems. In fact, by using signed networks, we can identify patterns and trends that reveal deeper insights into the dynamics of these complex systems, which may not be immediately visible through other analytical methods.
A central question in the analysis of signed networks is how well positive and negative edges align with Structural Balance Theory (SBT) (Cartwright and Harary, 1956). SBT asserts that positive relations should exhibit transitivity, as expressed by the saying “the friend of my friend is my friend,” while negative relations should demonstrate anti-transitivity, as in “the enemy of my enemy is my friend.”
To explore these dynamics, our approach utilises the exponential random graph modelling (ERGM) framework (Lusher, Koskinen and Robins, 2013), which is particularly suited to modelling the complex interplay between positive and negative edge configurations. ERGMs allow us to incorporate terms that capture relational tendencies, such as the clustering of positive edges or the formation of unexpected negative edges between individuals with similar nodal covariate information. By accounting for these relational patterns, we can accurately assess the statistical significance of balanced configurations and their impact on the overall structure of the network (Fritz et al., 2022).
In this paper, we examine undirected signed network panel data, which consists of cross-sectional snapshots of network relationships observed at regular intervals. In particular, we focus on the analysis of political relationships among U.S. Senators, measured across successive U.S. Congresses (Neal, 2020, 2014). Each snapshot captures the state of signed interactions within the Senate during each legislative year, offering insight into the evolving dynamics of political alliances and rivalries over time.
We extend the temporal exponential random graph model (TERGM) framework (Robins and Pattison, 2001; Hanneke, Fu and Xing, 2010) to better accommodate the dynamics of signed interactions in longitudinal network data. Specifically, we incorporate a separable specification (Krivitsky and Handcock, 2014) that distinguishes between the formation and persistence of edges, enabling a more detailed understanding of temporal dependence. Building on the decomposition introduced by Lerner (2016), we further enhance the TERGM by modelling the signed network as the outcome of two interrelated processes: an interaction process and a conditional sign process. This extension allows for a more flexible and interpretable parameterisation that captures not only how relationships form and persist over time, but also how structural balance – through both balanced and imbalanced signed configurations – shapes the evolution of network interactions.
The paper is structured as follows. In Sections 2 and 3, we provide an overview of ERGMs and temporal ERGMs respectively. Section 4 offers a comprehensive definition of signed networks and presents an overview of Structural Balance Theory (SBT). In Sections 5 and 6, we introduce a multi-layer interpretation of the signed process, distinguishing between the interaction process and the conditional sign process. We describe how these two components are incorporated into the ERGM generative framework to capture the dynamics of signed relationships. Section 7 combines temporal separability with layer separability to define separable 2-layer temporal ERGMs. Here, we define a Markov process based on binary ERGM terms, which allows for a flexible handling of Structural Balance Theory configurations. In Section 8, we outline a Bayesian inferential approach, which quantifies the uncertainty in the posterior model parameter estimates by using an adaptive Metropolis-Hastings approximate exchange algorithm (Haario, Saksman and Tamminen, 2001; Caimo and Friel, 2011) to facilitate sampling from the doubly-intractable posterior distribution. Finally, Section 9 demonstrates the utility of our modelling approach by applying it to the analysis of political relationships among U.S. Senators. We assess model fit using posterior predictive checks to evaluate the model performance. Section 10 concludes with a discussion on potential extensions and future directions.
2 Exponential random graph models
The relational structure of a network graph is described by a random adjacency matrix $\mathbf{y}$ of a graph on $n$ nodes (actors) and a set of edge variables such that $y_{ij}=1,$ if node $i$ is connected to node $j$ ($i\sim j$); $y_{ij}=0,$ if node $i$ is not connected to node $j$ ($i\not\sim j$).
Exponential random graph models (ERGMs) (Holland and Leinhardt, 1981; Strauss and Ikeda, 1990) are a particular class of discrete linear exponential families with probability mass function:
$p(\mathbf{y}\mid\boldsymbol{\varphi})=\exp\{\boldsymbol{\varphi}^{\top}\mathbf{s}(\mathbf{y})-\kappa(\boldsymbol{\varphi})\},$ (1)
where $\mathbf{s}(\mathbf{y})\in\mathbb{R}^{Q}_{>0}$ is a known vector of $Q$ sufficient statistics (Besag, 1974), $\boldsymbol{\varphi}\in\mathbb{R}^{Q}$ is the associated parameter vector, and $\kappa(\boldsymbol{\varphi})$ a normalising constant which is computationally intractable to evaluate for all but trivially small graphs. The dependence hypothesis at the basis of these models is that edges self organize into small structures called configurations. There is a wide range of possible network configurations (Robins et al., 2007) which give flexibility to adapt to different contexts. A positive (or negative) value of the parameter $\varphi$ indicates a tendency for the configuration represented by $s(\mathbf{y})$ to appear more (or less) frequently in the data than would be expected under an Erdős–Rényi random graph model with an edge probability of 0.5.
The ERGM likelihood defined in Equation (1) can be generalised by allowing the parameter to vary non-linearly within the exponential family, resulting in a curved exponential family (Hunter, 2007). Monte Carlo methods (Hunter and Handcock, 2006) can be used to estimate the decay parameter of the geometrically weighted network statistics introduced by Snijders et al. (2006) and currently used to alleviate ERGM degeneracy issues (Handcock, 2003).
3 Temporal exponential random graph models
Temporal exponential random graph models (TERGMs) (Robins and Pattison, 2001; Hanneke, Fu and Xing, 2010) describe the joint distribution of a dynamic network sequence by assuming:
$p(\mathbf{y}_{1},\dots,\mathbf{y}_{T}\mid\mathbf{y}_{0},\boldsymbol{\varphi})$ $=\prod_{t=1}^{T}p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\dots,\mathbf{y}_{0},\boldsymbol{\varphi}\right)$
$=\prod_{t=1}^{T}\exp\{\boldsymbol{\varphi}^{\top}\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1},\dots,\mathbf{y}_{0}\right)-\kappa(\boldsymbol{\varphi},\mathbf{y}_{t-1}\dots,\mathbf{y}_{0})\},$
where $\boldsymbol{\varphi}$ parametrises the influence of sufficient network statistics $\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1},\dots,\mathbf{y}_{0}\right)$ on the likelihood and $\kappa(\boldsymbol{\varphi},\mathbf{y}_{t-1},\dots,\mathbf{y}_{0})$ is a normalising constant.
Assuming first-order Markov dependence simplifies the model by limiting the dependence of $\mathbf{y}_{t}$ to $\mathbf{y}_{t-1}.$ This is of course computationally convenient and empirically justified in many settings:
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\dots,\mathbf{y}_{0},\boldsymbol{\varphi}\right)=p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\varphi}\right)\propto\exp\left\{\boldsymbol{\varphi}^{\top}\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1}\right)\right\}.$ (2)
The separable temporal parametrisation (STERGM) proposed by Krivitsky and Handcock (2014) provides a convenient control over incidence and duration of edges and separate interpretation between consecutive network observations. We define the formation network $\mathbf{y}^{\mathcal{F}}=\mathbf{y}_{t-1}\cup\mathbf{y}_{t}$; and the persistence network $\mathbf{y}^{\mathcal{P}}=\mathbf{y}_{t-1}\cap\mathbf{y}_{t}$; and assume $\mathbf{y}^{\mathcal{F}}$ and $\mathbf{y}^{\mathcal{P}}$ are conditionally independent given $\mathbf{y}_{t-1}$ so that:
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\varphi}^{\mathcal{F}},\boldsymbol{\varphi}^{\mathcal{P}}\right)=p\left(\mathbf{y}^{\mathcal{F}}\mid\mathbf{y}_{t-1},\boldsymbol{\varphi}^{\mathcal{F}}\right)\times p\left(\mathbf{y}^{\mathcal{P}}\mid\mathbf{y}_{t-1},\boldsymbol{\varphi}^{\mathcal{P}}\right),$ (3)
where $\boldsymbol{\varphi}^{\mathcal{F}}$ and $\boldsymbol{\varphi}^{\mathcal{P}}$ describe respectively the edge formation and persistence process from $t-1$ to $t.$ Then, given $\mathbf{y}^{\mathcal{F}},\mathbf{y}^{\mathcal{P}}$ and $\mathbf{y}_{t-1},$ we have $\mathbf{y}_{t}=\mathbf{y}^{\mathcal{P}}\cup\left(\mathbf{y}^{\mathcal{F}}\;\backslash\;\mathbf{y}_{t-1}\right).$
The STERGM framework addresses a key limitation of standard TERGMs by separating the modelling of edge formation and persistence. This separation allows for more interpretable and flexible modelling of dynamic networks, especially when edge formation and dissolution are governed by different mechanisms. For example, in a friendship network, a new friendship (formation) and the maintenance of an existing friendship (persistence) may be influenced by different factors — proximity may drive formation, while mutual friends may support persistence.
4 Signed network structures
A signed network consists of a set of $n$ nodes and signed edges between pairs of them. Its relational structure is described by a random adjacency matrix with elements the set of edge variables $y_{ij}\in\{+1,-1,0\}$ respectively indicating a positive, negative of no relation between node $i$ and node $j.$ As for binary networks, signed networks may be directed or undirected depending on the nature of the relationships between the nodes. In this paper, we focus on undirected signed networks, although the methodologies introduced can be easily extended to the directed case.
Structural balance theory (SBT) examines how positive and negative relationships between actors evolve and ultimately stabilise within a network. It focuses on the dynamics and eventual equilibrium of these relationships in a signed network (Cartwright and Harary, 1956). SBT is commonly illustrated using the familiar concepts of friendship and enmity. These terms provide a clear and intuitive understanding of the theory’s fundamental principles. Balanced triadic configurations include: (1) a friend of a friend is a friend, and (2) an enemy of a friend is an enemy. In contrast, unbalanced configurations are: (3) a friend of an enemy is an enemy, and (4) an enemy of an enemy is a friend. Configuration (1) corresponds to positive triangles (Figure 1 (a)), while configuration (2) corresponds to triangles with two negative edges (Figure 1 (b)). These two patterns are part of the so-called strong balance structure (Harary, 1953). Davis (1967) later introduced the weak balance structure, under which configurations comprising triangles with all negative edges (Figure 1 (c)) are also regarded as balanced. Conversely, triadic configurations with exactly two positive edges and one negative edge (Figure 1 (d)) are considered unbalanced.
Figure 1: Triadic configurations of a signed network. Solid lines represent positive interactions, dashed lines represent negative interactions.
The underlying motivation for structural balance is dynamic and based on how unbalanced triangles should resolve to balanced ones. This situation has led naturally to a search for a full dynamic theory of structural balance. Yet finding systems that reliably guide networks to balance has proved to be a challenge in itself (Marvel et al., 2011).
5 Multi-layer interpretation
Multi-layer ERGMs (Krivitsky, Koehly and Marcum, 2020) extend the traditional ERGM framework to model networks with multiple types of interactions between nodes, with each layer representing a distinct kind of relationship.
In the context of signed networks, we can decompose the signed edge process into the product of two network layers $\mathbf{y}=\mathbf{x}\mathbf{z},$ where $\mathbf{x},$ with dyadic elements $x_{ij}=0,$ if $i\sim j$; $x_{ij}=1,$ if $i\not\sim j,$ is a binary adjacency matrix encoding the presence of edges between nodes; and $\mathbf{z}$ with dyadic elements $z_{ij}=-1,$ if $i\overset{-}{\sim}j$; $z_{ij}=+1,$ if $i\overset{+}{\sim}j$ is a binary adjacency matrix encoding the sign of the edges between nodes. Clearly, this can be viewed as a special case of a multi-layer network, where the sign information is meaningful only in the presence of an actual interaction. In other words, the sign layer is effectively constrained by the interaction layer. This dependency will be further discussed in the following section.
However, one could also argue that signed relations may exist independently of observed interactions, i.e., the sign of a relationship could be latent, with the interaction merely serving to reveal or activate a pre-existing evaluative stance. In this view, the act of interaction exposes underlying sentiments such as trust, hostility, or alliance that already shape the relational landscape. This interpretation aligns, for example, with sociological theories of latent affective edges (Heider, 1946) and is particularly relevant in contexts where relationships are long-standing or embedded in broader social structures. In such settings, the absence of interaction does not necessarily imply neutrality or indifference, but rather a dormant state of the relationship.
We assume that the observed signed network follows a 2-layer ERGM with the probability mass function:
$p\left(\mathbf{y}\mid\boldsymbol{\vartheta}\right)=\exp\left\{\boldsymbol{\vartheta}^{\top}\mathbf{s}(\mathbf{x},\mathbf{z})-\kappa(\boldsymbol{\vartheta},\mathbf{x},\mathbf{z})\right\},$ (4)
where $\mathbf{s}(\mathbf{x},\mathbf{z})$ incorporates network statistics that model the two layers jointly. Figure 2 (a). The modelling approach for signed networks proposed by Fritz et al. (2022) is equivalent to a 2-layer ERGM which represents an ERGM with a uniform categorical reference measure.
6 Separable modelling
Lerner (2016) highlights the difference between marginal and conditional tests of structural balance, showing that other network effects affecting the likelihood of interaction can hide structural balance effects. This can make it seem as though balance theory is incorrect, even when nodes actually prefer balanced triangles over unbalanced ones. For this reason, to properly test balance theory, we should specify and estimate the conditional probability of an edge being positive or negative.
Unlike the joint model described in the previous section, this separable approach allows for conditional independent modelling of the structure of edges and their signs.
This has practical benefits: it simplifies model estimation, and it also aligns with situations where the factors driving connections (like shared interests or affiliations) are distinct from those affecting relationship quality (like competition or past conflict). For example, in a social network context, the presence of an edge between nodes might be explained by organisational requirements or physical proximity, while the sign could depend on trust, cooperation, or conflict.
The marginal probability of $\mathbf{y}$ can be decomposed into the product of the interaction process $\mathbf{x}$ and the conditional sign process given the presence of interaction $\mathbf{z}\mid\mathbf{x}$:
$p\left(\mathbf{y}\mid\boldsymbol{\zeta},\boldsymbol{\xi}\right)$ $=p\left(\mathbf{x}\mid\boldsymbol{\xi}\right)\times p\left(\mathbf{z}\mid\mathbf{x},\boldsymbol{\zeta}\right)$ (5)
$=\exp\left\{\boldsymbol{\xi}^{\top}\mathbf{s}\left(\mathbf{x}\right)-\kappa(\boldsymbol{\xi})\right\}\times\exp\left\{\boldsymbol{\zeta}^{\top}\mathbf{s}\left(\mathbf{z};\mathbf{x}\right)-\kappa(\boldsymbol{\zeta},\mathbf{x})\right\},$
where $\boldsymbol{\xi}$ governs the marginal interaction process, while $\boldsymbol{\zeta}$ governs the conditional sign process within the interaction structure. Figure 2 (b) provides a graphical representation of the model, illustrating that the two parameters, $\boldsymbol{\xi}$ and $\boldsymbol{\zeta}$, are conditionally independent given the interaction process $\mathbf{x}$. Consequently, we define Model (5) as a separable 2-layer ERGM.
Figure 2: Graphical representation of (a) 2-layer ERGMs defined in Equation (4) and (b) separable 2-layer ERGMs defined in Equation (5).
7 Separable 2-layer temporal models
The separable model defined in Equation (5), when combined with the separable temporal model family specified in Equation (3), results in a separable 2-layer TERGM, with the transition probability defined as:
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}}\right)=$ $\;p\left(\mathbf{x}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}}\right)\times p\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\right)$ (6)
$=$ $\;p\left(\mathbf{x}_{t}^{\mathcal{F}}\mid\mathbf{x}_{t-1},\boldsymbol{\xi}^{\mathcal{F}}\right)\times p\left(\mathbf{x}_{t}^{\mathcal{P}}\mid\mathbf{x}_{t-1},\boldsymbol{\xi}^{\mathcal{P}}\right)\times$
$\;p\left(\mathbf{z}_{t}^{\mathcal{F}}\mid\mathbf{x}_{t}^{\mathcal{F}},\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}}\right)\times p\left(\mathbf{z}_{t}^{\mathcal{P}}\mid\mathbf{x}_{t}^{\mathcal{P}},\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{P}}\right)$
$=$ $\exp\left\{{\boldsymbol{\xi}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\xi}^{\mathcal{F}}};\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\zeta}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\zeta}^{\mathcal{F}}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\xi}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\xi}^{\mathcal{P}}};\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\zeta}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\zeta}^{\mathcal{P}}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1})\right\},$
where $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{F}}$ describe the contribution of the interaction network effects and the edge sign network effects to the formation process, respectively; and where $\boldsymbol{\xi}^{\mathcal{P}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}$ describe the contribution of the interaction network effects and the edge sign network effects to the persistence process, respectively.
Model (6) inherits the separability property within each time interval from Model 3 in a way that $\boldsymbol{\xi}$ and $\boldsymbol{\zeta}$ are conditionally independent given $\mathbf{y}_{t-1}$ and $\mathbf{x}_{t}$ (and therefore $\mathbf{x}^{\mathcal{F}}_{t}$ and $\mathbf{x}^{\mathcal{P}}_{t}$) in each time interval. This means that while the formation and sign layers are modelled as conditionally independent within each time interval, Model (6) does not necessarily imply temporal independence across time. In fact, the structure of $\mathbf{x}_{t}$ may be influenced not only by $\mathbf{x}_{t-1}$, but also by $\mathbf{z}_{t-1}$, and similarly, $\mathbf{z}_{t}$ may depend not only on $\mathbf{z}_{t-1}$ but also on $\mathbf{x}_{t-1}$. In other words, the presence or absence of an interaction at time $t-1$ may explain not only the formation or dissolution of an edge at time $t$, but also changes in its sign; likewise, the sign structure at $t-1$ may influence not only whether sign persist or switch, but also whether new edges appear or existing ones disappear. This form of cross-layer temporal dependence is a key feature of the model, and reflects the fact that interaction and sign dynamics are often intertwined in practice. To achieve full separability over time between the two processes, we need to assume that at any time point $t,$ $\mathbf{x}^{\mathcal{F}}$ and $\mathbf{x}^{\mathcal{P}}$ (and therefore $\mathbf{x}$) do not depend on $\mathbf{z}_{t-1},$ i.e., at any time point $t,$ $\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)=\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{x}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)=\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{x}_{t-1}\right).$
Figure 3 provides a graphical representation of the separable 2-layer TERGM and illustrates that, given the interaction processes $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, the corresponding formation and persistence interaction processes – $\mathbf{x}_{1}^{\mathcal{F}}$, $\mathbf{x}_{2}^{\mathcal{F}}$, $\mathbf{x}_{1}^{\mathcal{P}}$, and $\mathbf{x}_{2}^{\mathcal{P}}$ – are fixed. This implies that $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{F}}$ are conditionally independent, as are $\boldsymbol{\xi}^{\mathcal{P}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}.$ However, temporal dependence between $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\xi}^{\mathcal{P}}$, as well as between $\boldsymbol{\zeta}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}$, still holds.
Figure 3: Graphical representation of separable 2-layer TERGMs for signed network data conditional on an initial network state (not included).
Table 1 presents the structure of all possible dyadic transitions in the signed network process over a given time interval, capturing how the relationships between node pairs evolve in terms of sign and connectivity.
Table 1: (a) Possible transitions of a single signed edge variable $Y_{ij}$ from $t-1$ to $t;$ (b) corresponding transitions according to the separable model defined in Equation (6).
(a) (b)
$Y_{i,j,t-1}$rightarrow $Y_{i,j,t}$ $Y_{i,j,t-1}$ $Y_{i,j}^{\mathcal{F}}$ $Y_{i,j}^{\mathcal{P}}$ $Y_{i,j,t}$
$X_{i,j,t-1}$ $Z_{i,j,t-1}$rightarrow $X_{i,j}^{\mathcal{F}}$ $Z_{i,j}^{\mathcal{F}}$ $X_{i,j}^{\mathcal{P}}$ $Z_{i,j}^{\mathcal{P}}$rightarrow $X_{i,j,t}$ $Z_{i,j,t}$
00 00 00 00 00 00
00 $+1$ 00 $1$ $+1$ 00 $1$ $+1$
00 $-1$ 00 $1$ $-1$ 00 $1$ $-1$
$+1$ 00 $1$ $+1$ $1$ $+1$ 00 00
$+1$ $+1$ $1$ $+1$ $1$ $+1$ $1$ $+1$ $1$ $+1$
$+1$ $-1$ $1$ $+1$ $1$ $+1$ $1$ $-1$ $1$ $-1$
$-1$ 00 $1$ $-1$ $1$ $-1$ 00 00
$-1$ $+1$ $1$ $-1$ $1$ $-1$ $1$ $+1$ $1$ $+1$
$-1$ $-1$ $1$ $-1$ $1$ $-1$ $1$ $-1$ $1$ $-1$
7.1 Model terms
As with any ERGM-based modeling framework, the selection of network statistics is highly flexible and context-dependent. Commonly used statistics, such as edges, homophily, and transitivity, are typically included in the model to capture key structural features. STERGMs are specified by two potentially different sets of network statistics for formation and persistence.
The network statistics $\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1}\right)$ include standard binary effects. These effects are based on the assumption that the layers $\mathbf{z}^{\mathcal{F}}$ and $\mathbf{z}^{\mathcal{P}}$ are endogenous, i.e., structures that are determined within the generative process itself, rather than being specified by external covariate information, while $\mathbf{x}^{\mathcal{F}}$, $\mathbf{x}^{\mathcal{P}}$, and $\mathbf{y}_{t-1}$ are treated as exogenous. The network statistics $\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)$ for the separable 2-layer TERGM (6) assume that the layers $\mathbf{x}^{\mathcal{F}}$ and $\mathbf{x}^{\mathcal{P}}$ are endogenous, conditional on $\mathbf{y}_{t-1}$.
Figure 4: Endogenous edgewise shared partners configurations: (a) esf+: positive edgewise shared friends; (b) ese+: positive edgewise shared enemies; (c) ese-: negative edgewise shared enemies; (d) esf-: negative edgewise shared friends.
As clearly demonstrated by Fritz et al. (2022), using lagged statistics that combine past and present edges can misrepresent the effects of SBT. For example, treating the presence of mutual friends or enemies at time $t-1$ as an exogenous covariate influencing the sign structure at time $t$ overlooks the fact that such covariates may not persist at time $t$, and may evolve in conjunction with other tie signs to form new balanced configurations. For this reason, in order to effectively model SBT effects, it is necessary to use endogenous network statistics that are implicitly dynamic – that is, they capture temporal dependencies without explicitly referencing past time points. It is important to emphasise that this reasoning also applies in our context, where SBT effects are estimated conditionally on the interaction structure. Indeed, in our case, assuming that such covariates remain constant would imply that both the interactions and their associated signs persist from time $t-1$ to time $t$, and are linked by an underlying structural dependency.
To test SBT effects, we adopt the geometrically-weighted statistics proposed by Fritz et al. (2022). These statistics can be included as binary terms $\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}}\right)$ and $\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}}\right)$ in separable 2-layer TERGMs (6) as the edgewise shared partner configurations shown in Figure 4 are used conditionally on the presence of interactions, resulting in binary configurations within a constrained support. For example, the geometrically weighted distribution of negative edgewise shared friends, $\texttt{gwesf}^{-}(\mathbf{y}_{t},\alpha)$, where $\alpha$ is the decay parameter, is equivalent to the geometrically weighted non-edgewise shared partners term, $\texttt{gwnsp}(\mathbf{z}_{t};\mathbf{x}_{t},\alpha)$, measured on $\mathbf{z}_{t}$ given $\mathbf{x}_{t}.$
8 Bayesian inference
We adopt a fully probabilistic Bayesian framework to estimate the separable 2-layer TERGM by defining a prior distribution over the model parameters $\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}})$, such that the posterior distribution is given by:
$\pi\left(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}},\boldsymbol{\xi}^{\mathcal{P}}\mid\mathbf{y}_{0:T}\right)$ $\propto\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}})\prod_{t=1}^{T}p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}}\right)$
$=\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}})\prod_{t=1}^{T}p\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\right)p\left(\mathbf{x}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}}\right).$
By assuming independent priors $\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}})=\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}})\times\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}})$ we preserve separability a posteriori:
$\pi\left(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{y}_{0:T}\right)=\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})\times\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}}\mid\mathbf{x}_{0:T}).$
The two posterior terms are doubly intractable since neither the corresponding likelihoods nor the marginal likelihoods are available. However, we can estimate both by adapting the approximate exchange algorithm (AEA) as proposed by Caimo and Friel (2011). Algorithm 1 outlines the AEA algorithm for posterior sampling of parameters $\boldsymbol{\zeta}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}$ given observations $\mathbf{z}_{0:T}$ and $\mathbf{x}_{0:T}$. Sampling from $\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}}\mid\mathbf{x}_{0:T})$ follows a similar procedure, but omits the dependence on signed edge states $\mathbf{z}_{t}$ during proposal generation for interaction structures $\mathbf{x}^{\mathcal{F}^{\prime}}$ and $\mathbf{x}^{\mathcal{P}^{\prime}}$.
Algorithm 1 AEA sampling from $\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})$
  Initialise $\boldsymbol{\zeta}^{\mathcal{F}(0)},\boldsymbol{\zeta}^{\mathcal{P}(0)}.$
  Compute $\mathbf{x}_{t}^{\mathcal{P}}$ and $\mathbf{x}_{t}^{\mathcal{P}},\forall t=1,\dots,T.$
  for $i=1,\dots,I$ do
     $\mathbf{i.}$ Propose $\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\sim h(\cdot)$
     for $t=1,\dots,T$ do
        a. Simulate $\mathbf{z}^{\mathcal{F}^{\prime}}\sim p\left(\cdot\mid\mathbf{x}_{t}^{\mathcal{F}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}}\right)$
        b. Simulate $\mathbf{z}^{\mathcal{P}^{\prime}}\sim p\left(\cdot\mid\mathbf{x}_{t}^{\mathcal{P}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{P}}\right)$
        c. Set $\mathbf{z}_{t}^{\prime}=\mathbf{z}_{t}^{\mathcal{P}^{\prime}}\cup\left(\mathbf{z}_{t}^{\mathcal{F}^{\prime}}\;\backslash\;\mathbf{z}_{t}\right)$
     end for
     $\mathbf{ii.}$ Set $(\boldsymbol{\zeta}^{\mathcal{F}(i)},\boldsymbol{\zeta}^{\mathcal{P}(i)})=(\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}})$ with probability:
$\min\left(1,\prod_{t=1}^{T}\frac{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\times\frac{\pi\left(\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\pi\left(\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\right).$ (7)
  end for
To address the sampling challenges arising from highly correlated parameters—including, as anticipated, correlations between the formation and persistence parameter groups—we employed an adaptive Metropolis–Hastings (AMH) proposal scheme (Haario, Saksman and Tamminen, 2001). Although the adaptive direction sampling scheme (used by default in Caimo and Friel (2011)) necessitates multiple chains and is less efficient over the long run, it remains valuable for sampling initial parameter values during the burn-in period.
The approximate likelihood ratio in (7) is computed as follows:
$\prod_{t=1}^{T}$ $\frac{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}$
$\hskip 10.81218pt=\exp\left\{\left(\boldsymbol{\zeta}^{\mathcal{F}(i-1)}-\boldsymbol{\zeta}^{\mathcal{F}^{\prime}}\right)^{\top}\sum_{t=1}^{T}\left[\mathbf{s}\left(\mathbf{z}^{\mathcal{F}^{\prime}}_{t};\mathbf{x}^{\mathcal{F}}_{t},\mathbf{y}_{t-1}\right)-\mathbf{s}\left(\mathbf{z}^{\mathcal{F}}_{t};\mathbf{x}^{\mathcal{F}}_{t},\mathbf{y}_{t-1}\right)\right]\right.$
$\hskip 28.45274pt\quad\left.+\left(\boldsymbol{\zeta}^{\mathcal{P}(i-1)}-\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)^{\top}\sum_{t=1}^{T}\left[\mathbf{s}\left(\mathbf{z}^{\mathcal{P}^{\prime}}_{t};\mathbf{x}^{\mathcal{P}}_{t},\mathbf{y}_{t-1}\right)-\mathbf{s}\left(\mathbf{z}^{\mathcal{P}}_{t};\mathbf{x}^{\mathcal{P}}_{t},\mathbf{y}_{t-1}\right)\right]\right\}.$
A key advantage of this approach is that network sampling from the likelihood (step a and b in Algorithm 1) can be conducted with significantly fewer auxiliary iterations and lower computational cost, thanks to the sampling space constraint and reduced complexity of the statistics which are binary.
8.1 Software
The new B2Lstergm package for R (R Development Core Team, 2025), developed alongside this paper, represents a significant extension of existing tools for Bayesian inference on temporal network models. Specifically designed for Bayesian separable 2-layer TERGMs on signed networks, B2Lstergm is currently under active development and aims to offer a more flexible and computationally efficient framework than what is currently available. It builds on and extends the modelling and computational capabilities of the Bergm package (Caimo et al., 2022, 2025), which implements Bayesian inference for static ERGMs. Additionally, the multi.ergm package (Krivitsky, 2025), which supports multi-layer ERGMs (Krivitsky, Koehly and Marcum, 2020), is employed here to simulate network statistics corresponding to the interaction and conditional sign processes of the two separable likelihood components described in Equation 6. This simulation relies on the efficient MCMC algorithms implemented in the ergm package (Hunter et al., 2008; Krivitsky et al., 2023; Handcock et al., 2025). Notably, although the stergm() function within the tergm package (Krivitsky and Handcock, 2025) typically relies on stochastic maximum likelihood estimation, it can alternatively use contrastive divergence when computational efficiency is a priority or convergence issues arise.
9 Application to US Congressional data
We employ our model to examine the evolution of political cooperation and opposition in the United States Congress using the signed network data inferred by Neal (2020, 2014) from bill co-sponsorship data via stochastic degree sequence model.
Our focus is on the 99th-101st Congresses (1985-1991), characterised by a hung government with a split majority. This period saw a republican president, Ronal Reagan, governing alongside a Democratic majority in the Senate. An examination of the Senate’s behaviour during this time can provide valuable insights into the dynamics of the Senate and its role in shaping policy during a period of significant upheaval (Rohde, 1991). This period encompasses several pivotal events, including the collapse of the Soviet Union (1989-1991), the Gulf War (1990-1991), and the Savings and Loan Crisis (1989-1991).
We concentrate our analysis on the 70 Senators who remained in office throughout the 99th-101st Congresses, in order to examine the behaviour of the network effects related to their party affiliation and structural balance.
Figure 5: US Senators signed network. Blue and red nodes represent Democratic and Republican Senators, respectively. Graphs: (a) signed relationships during the 99th Congress (initial state); (b) signed relationships during the 100th Congress; (c) signed relationships during the 101th Congress.
The observed signed networks displayed in Figure 5 exhibit considerable interaction density. Specifically, the formation interaction density rises from just below 0.60 to nearly 0.61, while the persistence interaction density increases from 0.26 to almost 0.31. Most interactions are negative, even within parties, across both the formation and persistence processes.
9.1 Model specification
We adopt a time-homogeneous separable 2-layer TERGM where the formation and persistence configurations are the same and the list of statistics used is specified below. Obviously, in the case of excessive non-stationarity across multiple network observations, the use of a homogeneous model could prove problematic.
As mentioned in the previous section, our focus is to assess the extent of the importance of weak balance structures, and thus we include (non-degenerate) configurations corresponding to the weakly balanced triads (Figure 1) in both the formation and persistence processes.
The interaction network exhibits such high density that it lacks discernible structural heterogeneity. Consequently, we restrict our analysis to the conditional signed process and the posterior distribution of its associated parameters, $\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})$. We define a separable 2-layer TERGM using the following network statistics:
  • 1.
    number of positive edges, ${}^{+},$ measuring the density of positive interactions;
  • 2.
[…]
  • 3.
[…]
  • 4.
[…]
  • 5.
[…]
  • 6.
[…]
We set the decay parameter $\alpha$ to $0.2$ for the geometrically-weighted degree statistic, and to $0.6$ for all geometrically-weighted edgewise shared partner-based statistics (Snijders et al., 2006).
We assign independent, weakly informative normal priors to the model parameters, informed in part by the structural characteristics of the initial network $\mathbf{y}_{0}$. Specifically, we set the prior means for both $\zeta^{\mathcal{F}}_{\text{edges}^{+}}$ and $\zeta^{\mathcal{P}}_{\text{edges}^{+}}$ to –1, and the prior mean for $\zeta^{\mathcal{F}}_{\text{gwesf}^{+}}(\alpha)$ is set accordingly. All prior variances are fixed at 25.
9.2 Posterior inference
Posterior inference is conducted using the adaptive Metropolis–Hastings approximate exchange algorithm (MH-AEA), as outlined in Algorithm (1). Each iteration employs $5000$ auxiliary draws for the network simulation, and the main Markov chain comprises $35000$ iterations, with the initial $10000$ discarded as burn-in. The MCMC trace plots (Figure 6) of simulated posterior parameters indicate that the MCMC chain is convergent to the stationary posterior distribution. Table 2 presents a summary of the marginal posterior parameter density estimates.
Figure 6: MCMC histograms and trace plots for each marginal posterior parameter distribution.
Posterior predictive checks are performed to evaluate how well the model captures the observed data. Simulated network statistics were generated from the posterior predictive distribution and compared to the observed statistics using boxplots (Figure 7). In these plots, the medians of the posterior predictive statistics are close to $0,$ indicating that the observed data are consistent with what the model predicts. This suggests that the model provides a good fit to the data in terms of model statistics.
Figure 7: Posterior predictive boxplots of centred cumulative differences over time between simulated and observed network statistics: $\sum_{t=1}^{T}\mathbf{s}_{q}\left(\mathbf{z}^{\mathcal{U}^{\prime}}_{t};\mathbf{x}^{\mathcal{U}}_{t},\mathbf{y}_{t-1},\zeta_{q}^{\mathcal{U}}\right)-\mathbf{s}\left(\mathbf{z}^{\mathcal{U}}_{t};\mathbf{x}^{\mathcal{U}}_{t},\mathbf{y}_{t-1},\zeta_{k}^{\mathcal{U}}\right),\mathcal{U}=\{{\mathcal{F},\mathcal{P}}\},q=1,\cdots,Q.$
Table 2: Summary of the posterior parameter density estimates.
Formation ($\mathcal{F}$) Persistence ($\mathcal{P}$)
Parameters Mean 0.025 0.5 0.975 Mean 0.025 0.5 0.975
$\zeta_{\texttt{edges}^{+}}$3.95 -5.81 -3.98 -2.00 -4.77 -5.90 -4.76 -3.68
$\zeta_{\texttt{homophily}^{+}(\texttt{rep})}$ 0.53 0.02 0.53 1.06 -0.28 -0.83 -0.27 0.25
$\zeta_{\texttt{gwdegree}^{+}(\alpha=0.2)}$0.29 -3.11 -0.31 2.49 0.81 -0.28 0.79 1.87
$\zeta_{\texttt{gwesf}^{+}(\alpha=0.6)}$ 2.42 1.46 2.44 3.38 1.56 1.05 1.56 2.11
$\zeta_{\texttt{gwese}^{+}(\alpha=0.6)}$1.41 -2.11 -1.40 -0.70 -0.58 -0.96 -0.58 -0.20
$\zeta_{\texttt{gwese}^{-}(\alpha=0.6)}$1.05 -1.41 -1.03 -0.75 -0.62 -0.81 -0.61 -0.48
Although the analysis is restricted to the subset of U.S. Senators who served continuously from the 99th to the 101st Congress, the posterior parameter estimates, presented in Table 2, yield several key insights into the dynamics of political interactions:
  • The strongly negative 95% posterior credible intervals for $\zeta_{\texttt{edges}^{+}}$ in both the formation and persistence processes suggest that the presence of an existing positive edge reduces the likelihood of forming or maintaining additional positive interactions. This may reflect either competition for alliances or strategic repositioning in response to institutional instability.
  • The 95% posterior credible interval for $\zeta_{\texttt{homophily}^{+}(\texttt{rep})}$ is slightly positive for formation but not significantly different from 0 for persistence, indicating that while Republican senators exhibit a modest tendency to form positive edges within the party, these interactions are not particularly stable over time. This instability may point to underlying intra-party conflict.
  • The estimate for $\zeta_{\texttt{gwdegree}^{+}(\alpha=0.2)}$ in both processes is not significantly different from zero, suggesting a relatively even distribution of interactions, implying no strong centralisation in the network structure.
  • The consistently positive 95% posterior credible intervals for $\zeta_{\texttt{gwesf}^{+}(\alpha=0.6)}$ provide robust support for the role of balanced triadic closure in both initiating and sustaining positive interactions. This is consistent with the theoretical expectation of structural balance and indicates a preference for cohesive, reinforcing social triangles.
  • The negative posterior estimates for $\zeta_{\texttt{gwese}^{+}(\alpha=0.6)}$ and $\zeta_{\texttt{gwese}^{-}(\alpha=0.6)}$ (also negative across both processes) suggest that unbalanced triadic structures – whether involving positive or negative edges – are systematically avoided. This aligns with strong structural balance theory.
Collectively, these findings suggest a polarised but strategically coherent interaction network, in which alliances remain fluid and structural balance emerges as a more stable organising mechanism than party affiliation.
10 Discussion
This paper introduces a novel separable temporal generative framework for modelling the dynamics of signed networks. By distinguishing between interaction effects and conditional sign processes, our approach preserves the flexibility and interpretability of traditional binary exponential random graph models while remaining consistent with the assumptions of structural balance theory (SBT). To empirically test SBT, we specifically incorporate endogenous effects that capture patterns predicted by the theory. Nonetheless, the framework remains fully compatible with the broader class of binary statistics commonly used in separable temporal ERGM (STERGM) processes, including exogenous and lagged covariates. This separation allows for a deeper understanding of how network relations evolve over time, making it suitable for complex systems in which the interplay of both positive and negative relationships is critical to the system’s overall dynamics.
We employed a fully probabilistic Bayesian approach to infer model parameters, utilising an adaptive Metropolis-Hastings approximate exchange algorithm to efficiently estimate the doubly intractable posterior distribution. This method enables us to account for the complexity of signed network structures and provides robust uncertainty estimates despite the computational challenges.
We analysed the political relationships among U.S. Senators during Ronald Reagan’s second term (1985–1989), uncovering the evolving dynamics of political alliances and rivalries within the Senate. By applying our framework with a focus on endogenous structural effects grounded in SBT, we identified key patterns of both supportive and antagonistic alliances, shedding light on the structural mechanisms underpinning shifts in political coalitions.
The flexibility and interpretability of our model open up several avenues for future research. One promising direction involves extending our approach to a continuous-time framework, such as those based on longitudinal ERGMs (Koskinen and Lomi, 2013; Koskinen, Caimo and Lomi, 2015), which could significantly enhance the model’s applicability to real-time, dynamic network data. This extension may involve the development of data augmentation strategies in which MCMC algorithms alternate between sampling from the conditional posterior distribution of the model parameters and generating plausible interaction paths that connect consecutive network snapshots. Another direction concerns the adaptation of the methodology to dynamic weighted signed networks, where the transition probabilities of the signed process follow a partially separable temporal model, as recently proposed by Kei, Chen and Padilla (2023). Crucially, the feasibility of these extensions will depend on the development of more efficient and scalable estimation algorithms. As network data grow in complexity and temporal resolution, computational bottlenecks become increasingly significant. Advancing the underlying inferential algorithm – through improved MCMC schemes, variational approximations (Tan and Friel, 2020) – will be essential to ensure the practical usability of these models in applied settings.